You recieve a box with books: 2 romances 3 mystery novels and four science fiction
A) if you select four books at random in succession with out replacement what is the probability that the fourth book is a romance?
B) if you select two books at random in succession without replacement what is the probability that the first book is a mystery novel given that the second book is a science fiction novel?
Answer #1

A) chances of 4th book being romance = (7*6*5*2) + (8*7*6*1) = 420+ 336=756

total = 9*8*7*6=3024

prob = 756/3024

B)Prob that 1st is mystery is 3/9.